OpenStudy (anonymous):

You plan to buy a $12,000 sailboat in four years. How much should you invest now, at 7.3% simple interest, to have enough for the boat in four years? (Round to the nearest cent.)

hero (hero):

You use the same formula, except you find P

hero (hero):

A = P(1+rt) 12,000 = P(1+.073*4) 12,000/(1.073*4) = P

OpenStudy (anonymous):

Thank you very much!

hero (hero):

wait, I think you need to multiply rt before you add the one

hero (hero):

12000/1.892 = P

hero (hero):

6342.49 = P

OpenStudy (anonymous):

Thank you! I appreciate this! This section of math is really confusing me because there are so many formulas. I have a few more questions I'm going to ask, could you please take a look at them? I'm trying to do them, but my answers have been off.

OpenStudy (lalaly):

12000/1.292

OpenStudy (anonymous):

Thank you!

hero (hero):

ugh

hero (hero):

I put 8 instead of 2

OpenStudy (anonymous):

oh no :(

hero (hero):

12000/1.892 = P should be 12000/1.292

hero (hero):

lalady backs me up all the time

OpenStudy (lalaly):

:D:D:D

hero (hero):

Yeah lalaly, you're the man!

hero (hero):

oh wait...

OpenStudy (anonymous):

haha I'm about to post another question if you guys want to help :)

OpenStudy (lalaly):

no hero you are the man

hero (hero):

:D

OpenStudy (anonymous):

Use the formula for simple interest, I = P × r × t, to find the missing quantity. I = $960; P = $6,000; r = 8%

hero (hero):

This is just plug and play

hero (hero):

We normally write it as I = PRT

OpenStudy (anonymous):

so it would be 2?

hero (hero):

I = PRT 960 = 6000 x .08 x T 960 = 480 x T 960/480 = T 2 = T Correct

OpenStudy (anonymous):

okay, thank you!

OpenStudy (anonymous):

The principal P is borrowed at a simple interest rate r for a period of time t. Find the loan's future value A, or the total amount due at time t. Round answer to the nearest cent. P = $38,000, r = 5.0%, t = 3 years Is this kind of the same deal as the ones before?

hero (hero):

Wow, you're still here?

hero (hero):

It's the same as the other one, except you have to find I I = PRT Find variable I

OpenStudy (anonymous):

Thank you!